OpenStudy (anonymous):

The line that is the perpendicular bisector of the segment whose endpoints are R(-1, 6) and S(5, 5)

OpenStudy (shamim):



OpenStudy (shamim):

let the mid point is T(x,y)

OpenStudy (shamim):

so\[x=\frac{ -1+5 }{ 2 }=2\]

OpenStudy (shamim):

and\[y=\frac{ 6+5 }{ 2 }=\frac{ 11 }{ 2 }\]

OpenStudy (shamim):

did u get it

OpenStudy (anonymous):

i need it in standard from

OpenStudy (anonymous):

i can get all that i just need help getting it in standard form

OpenStudy (shamim):

now the slope of ur given line RS\[m _{1}=\frac{ 6-5 }{ -1-5 }=-\frac{ 1 }{ 6 }\]

OpenStudy (shamim):

ok

OpenStudy (shamim):

\[y-\frac{ 11 }{ 2 }=6(x-2)\]

OpenStudy (shamim):

is it wt u want

OpenStudy (shamim):

u can also write \[12x-2y=13\]

OpenStudy (shamim):

m i wrong

OpenStudy (anonymous):

no your right thank you.

OpenStudy (shamim):

welcome